For the population of 2,5,4,3,2



Construct the sampling distribution of the mean x of size 3 without replacement

1
Expert's answer
2022-01-05T05:19:31-0500

Mean


"\mu=\dfrac{2+5+3+4+2}{5}=3.2"

Variance


"\sigma^2=\dfrac{1}{5}\big((2-3.2)^2+(5-3.2)^2+(3-3.2)^2""+(4-3.2)^2+(2-3.2)^2\big)=1.36"


Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{1.36}\approx1.1662"



We have population values "2,5,3,4,2" population size "N=5" and sample size "n=3." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{3}=10""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,2,3 & 7/3 \\
 \hdashline
 2 & 2,2,4 & 8/3 \\
 \hdashline
 3 & 2,2,5 & 3 \\
 \hdashline
 4 & 2,3,4 & 3 \\
 \hdashline
 5 & 2,3,5 & 10/3 \\
 \hdashline
 6 & 2,4,5 & 11/3 \\
 \hdashline
 7 & 2,3,4 & 3 \\
 \hdashline
 8 & 2,3,5 & 10/3 \\
 \hdashline
 9 & 2,4,5 & 11/3 \\
 \hdashline
 10 & 3,4,5 & 4 \\ \hline
\end{array}"



The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 7/3 & 1& 1/10 & 7/30 & 49/90 \\
 \hdashline
 8/3 & 1& 1/10 & 8/30 & 64/90 \\
 \hdashline
 3 & 3 & 3/10 & 27/30 & 243/90 \\
 \hdashline
 10/3 & 2 & 2/10 & 20/30 & 200/90\\
 \hdashline
 11/3 & 2 & 2/10 & 22/30 & 242/90 \\
 \hdashline
 4 & 1 & 1/10 & 12/30 & 144/90 \\
 \hdashline
 Total & 10 & 1 & 96/30 & 942/90 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{96}{30}=3.2"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=3.2=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2""=\dfrac{942}{90}-(\dfrac{96}{30})^2=\dfrac{68}{300}=\dfrac{17}{75}""\sqrt{Var(\bar{X})}=\sqrt{\dfrac{17}{75}}\approx0.4761"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{1.36}{3}(\dfrac{5-3}{5-1})""=\dfrac{0.68}{3}=\dfrac{17}{75}, True"